THE WEAKLY–BOUND CO_{2}–ACETYLENE COMPLEX: FUNDAMENTAL AND TORSIONAL COMBINA-TION BAND IN THE CO_{2} ν_{3} REGION

C. LAUZIN, Laboratoire de Chimie quantique et Photophysique, CP160/09 Faculté des Sciences, Université Libre de Bruxelles (U.L.B.), Ave. Roosevelt, 50 B-1050 Brussels, Belgium; \underline{J. NOROOZ OLIAEE}, M. REZAEI, N. MOAZZEN-AHMADI, Department of Physics and Astronomy, University of Calgary, Calgary, AB T2N 1N4, Canada.

Infrared spectrum of the weakly-bound CO_{2}–C_{2}H_{2} complex in the region of the CO_{2} ν_{3} fundamental band (∼ 2349 cm^{−1}) is observed in a pulsed supersonic slit jet expansion using a tunable diode laser probe. Two bands are observed and analyzed: the fundamental (C–O asymmetric stretch) and a combination involving the intermolecular torsional (out-of-plane bend) vibration. The resulting torsional frequency is 44.385(10) cm^{−1}. This represents the first observation of an intermolecular frequency for carbon dioxide-acetylene complex. A comparison between the results obtained here and those previously reported for N_{2}O–C_{2}H_{2} complex ^{a} is discussed.

aM. Dehghany, Mahin Afshari, J. Norooz Oliaee, N. Moazzen-Ahmadi, A. R. W. McKellar, Chem. Phys. Lett. 473 (2009), 26-29.